salary data was collected from each school system within a state. it was found that the mean annual teacher salary was normally distributed. the mean annual teacher salary for all of the school systems was $35,000 and the standard deviation was $19,500. school system a has a mean annual teacher salary of $47,500. what percent of the school systems pay a lower mean annual teacher salary than school a?
Answer by stanbon(75887)

salary data was collected from each school system within a state. it was found that the mean annual teacher salary was normally distributed. the mean annual teacher salary for all of the school systems was $35,000 and the standard deviation was $19,500. school system a has a mean annual teacher salary of $47,500. what percent of the school systems pay a lower mean annual teacher salary than school a?
-----------------
z(47500) = (47500-35000)/19500 = 0.64
------
P(x < 47500) = P(z < 0.64) = normalcdf(-100,0.64) = 0.7389
------
Cheers,
Stan H.
------------